# STD-LLM: Understanding Both Spatial and Temporal Properties of Spatial-Temporal Data with LLMs

### v1

@@ 4.1 Datasets
The experiments conducted on four publicly available real-world traffic datasets (PEMS03, PEMS04, PEMS07, PEMS08) to validate the effectiveness and generalization capability of STD-LLM. PEMS means Caltrans Performance Measurement System (PeMS)[44], which is a system that collects data from the California highway network. The details of the four datasets are provided in Appendix C.2.
To test the model’s imputation capabilities under different scenarios, we generated missing data for testing using two type of missing C.3: RM (random missing), CM (spatial-temporal continuity missing) and two rates of missing - 30%, 70% on the dataset. Additionally, following the self-supervised training approach of CSDI[42], we generated condition missing for training. Condition missing refers to the artificially created missing data, used to generate training samples.
@@ 4.2 Experimental Settings
To comprehensively compare the performance of models, we selected five representative baselines for the forecasting task, including Long Short-Term Memory network(LSTM)[45], ASTGCN[30], AGCRN[3], ASTGNN[4], PDFormer[31], OFA[12], and STGLLM[13]. For the imputation task, we also chose four well-known baselines, which are Brits[5], E2GAN[36], mTAN[40], and PriSTI[6].
The experiments were conducted on an Ubuntu 23.04 operating system with a PyTorch version of 2.1.2. The AdaW optimizer was utilized to optimize the model. The learning rate was set to $1\times 10^{-3}$. The training epoch is 500 with an early stopping mechanism based on a patience of 50 epochs. The loss function used was L1 loss. We use GPT-2[46] as the pre-trained LLM and utilized only its initial three layers. STD-LLM employs the hyperparameters identified through tuning on the PEMS08 validation set for all datasets, ensuring the model’s generalization capability across different datasets. The detail of experiments settings are depicted in Appendix C.
@@ 4.3 Overall Performance
We measure the model’s performance using three widely used metrics for regression tasks: Mean Absolute Error (MAE), Root Mean Square Error (RMSE), and Mean Absolute Percentage Error (MAPE). Table 1 shows the forecasting performances comparison between our proposed STD-LLM and baselines. Bold indicates the best results, while underline denotes the second-best results. From Table 1, we can observe that: 1) The accuracy of the time series models such as LSTM and OFA is not good enough compared to methods that can handle spatial-temporal correlations in data. This indicates that relying solely on temporal correlation is insufficient to achieve high-precision forecasting; 2) The results obtained from OFA and STGLLM suggest that it is challenging for LLM to capture key information such as topological structure and higher-order spatiotemporal dependencies relying solely on its capabilities. This implies that it is necessary for us to design spatial and temporal tokenizers, topology-aware node embeddings, and hypergraph learning module to assist the LLM in understanding spatial and temporal properties. 3) Our proposed method achieve the best or second-best results on the PEMS04, PEMS07, and PEMS08 datasets.
Table 2 presents the imputation performances comparison between our proposed STD-LLM and baselines. Based on the comparison, we can observe that: 1) Our method achieves state-of-the-art performance in most scenarios. Although it underperforms in the CM 30%, it still attains the second-best results; 2) As the rate of missing data increases, the performance advantage of STD-LLM becomes more pronounced. This suggests that our method can more accurately capture higher-order correlations and other critical information from a small amount of data, demonstrating the capability to handle highly incomplete data; 3) STD-LLM exhibits a significant advantage in the RM case, particularly with a substantial improvement over other baseline models in the RM 70% case.
In summary, the experimental results demonstrate that STD-LLM has strong performance and can achieve high accuracy in both forecasting and imputation tasks.
Table 1: Forecasting performance.
Model PEMS03 PEMS04 PEMS07 PEMS08
MAE RMSE MAPE(%) MAE RMSE MAPE(%) MAE RMSE MAPE(%) MAE RMSE MAPE(%)
LSTM 20.62 33.54 28.94 26.81 40.74 22.33 29.71 45.32 14.14 22.19 33.59 18.74
ASTGCN 17.85 29.88 17.65 22.42 34.75 15.87 25.98 39.65 11.84 18.86 28.55 12.50
AGCRN 15.98 28.25 15.23 19.83 32.26 12.97 22.37 36.55 9.12 15.95 25.22 10.09
ASTGNN 14.78 25.00 14.79 18.60 30.91 12.36 20.62 34.00 8.86 15.00 24.70 9.50
PDFormer 14.74 25.59 15.35 18.31 29.97 12.10 19.83 32.87 8.53 13.58 23.51 9.05
OFA 20.96 33.43 19.11 27.37 42.99 17.97 30.53 47.51 12.98 21.89 34.63 13.30
STGLLM 15.26 24.11 15.73 20.00 32.11 13.69 21.98 35.02 9.72 15.53 24.74 10.15
STD-LLM (ours) 14.89 25.38 15.45 18.31 30.32 11.96 19.71 32.99 8.32 13.85 23.39 9.04
Table 2: Imputation performance on PEMS08. RM means the random missing. CM means the spatial-temporal continuity missing.
Model RM 30% RM 70% CM 30% CM 70%
MAE RMSE MAPE(%) MAE RMSE MAPE(%) MAE RMSE MAPE(%) MAE RMSE MAPE(%)
BRITS 21.28 29.78 7.57 29.56 41.52 10.79 33.65 46.75 11.97 38.33 53.60 13.90
E2GAN 23.75 37.38 15.35 27.55 41.99 17.52 25.66 40.55 16.44 31.58 50.24 19.17
mTAN 20.29 32.62 12.71 21.23 33.98 12.89 24.50 40.23 15.07 31.42 49.10 18.96
PriSTI 14.67 23.17 9.55 25.54 36 16.73 17.8 34.67 10.71 22.93 40.9 14.43
STD-LLM (ours) 12.14 19.44 8.06 14.38 22.92 9.46 19.36 34.53 12.30 22.18 39.26 13.56
@@ 4.4 Few-shot and Zero-shot Performance
To extensively evaluate the performance of LLMs on spatial-temporal tasks, we conduct few-shot and zero-shot experiments. The few-shot experiments evaluate the forecasting performance of STD-LLM on the PEMS04 and PEMS08 datasets using only the first 5%, 10%, and 20% of the training samples, with results summarized in Table 4. The zero-shot experiments evaluate the forecasting performance of the pre-trained model when applied to other datasets, with the results compiled in Table 4.
From Table 4, it can be observed that we only require 5% of the training samples to achieve performance comparable to that of LSTM trained on the full dataset. When the training samples are increased to 20%, the performance surpasses that of ASTGCN. This indicates that STD-LLM has excellent few-shot learning capabilities, which supports its application in scenarios where data is scarce. The results in Table 4 demonstrate that STD-LLM can exhibit acceptable performance when directly transferred to datasets with different temporal scopes and graph structures from the training set, without undergoing any training. Further experiments are presented in the Appendix F.
@@
MAE RMSE MAPE(%) MAE RMSE MAPE(%)
5% 26.37 40.98 18.80 20.71 32.22 14.51
10% 24.15 37.91 16.57 19.60 31.09 12.85
20% 21.43 34.83 14.17 16.88 27.76 10.84
@@
MAE RMSE MAPE(%)
$\xrightarrow{}$08 25.95 39.60 21.74
$\xrightarrow{}$04 29.29 46.01 22.46
$\xrightarrow{}$03 23.58 35.50 38.27
$\xrightarrow{}$07 33.79 51.63 18.73
@@ 4.5 Ablation Study
Table 5: Ablation study.
Method PEMS04 PEMS08 PEMS08 RM 70% PEMS08 CM 70%
MAE RMSE MAPE(%) MAE RMSE MAPE(%) MAE RMSE MAPE(%) MAE RMSE MAPE(%)
w/o Temporal Token 18.57 30.61 12.26 13.92 23.47 9.11 14.81 23.40 10.11 24.71 41.24 17.02
w/o Virtual Node 18.73 30.86 12.47 14.10 23.76 9.23 14.67 23.35 9.71 23.18 40.01 14.71
w/o Hypergraph 18.54 30.73 12.11 14.31 24.18 9.21 14.59 23.15 9.88 25.64 42.65 17.18
w/o LLM 18.96 30.88 12.82 14.25 23.47 9.36 21.66 36.60 21.12 104.30 133.72 197.80
STD-LLM 18.31 30.32 11.96 13.85 23.39 9.04 14.38 22.92 9.46 22.18 39.26 13.56
We implement ablation study to evaluate the effectiveness of each component. The ablation experiments compared our proposed model with the following variants: 1) w/o Temporal Token, removing the temporal tokens in Equation 6; 2) w/o Virtual Node, removing the virtual node; 3) w/o Hypergraph, removing the Hypergraph Learning module in Equation 7; 4) w/o LLM, replacing the LLM with a transformer encoder of the same layer and hidden dimension.
We summerize the results of the ablation study in Table 5. Based on the results, we can identify the following: 1) Temporal Token and Virtual Node can enhance the model’s forecasting and imputation performance with a relatively small computational cost. 2) The hypergraph module not only reduces the number of tokens, thereby increasing efficiency, but also significantly boosts the model’s performance. 3) The universal pattern recognition capabilities embodied in the pre-trained parameters of the LLM are useful for both spatiotemporal forecasting and imputation. Thus, LLM indeed play a substantial role in spatial-temporal forecasting and imputation, especially in scenarios with complex spatial-temporal dependencies, like the imputation task on PEMS08 CM 70%.
@@ 5 Conclusion
In this paper, we propose STD-LLM, a unified framework for spatial-temporal forecasting and imputation based on LLMs. Through our explicitly designed spatial and temporal tokenizers, STD-LLM can effectively understand both spatial and temporal properties of spatial-temporal data. To enhance the LLM’s understanding and utilization of the topology structure in spatial-temporal data, we have designed topology-aware node embeddings. Considering the impact of external systems, we have also introduced virtual nodes. In addition, we design a hypergraph learning module to enables the LLM to capture the non-pairwise and higher-order correlations by constructing region-level tokens. Extensive experiments demonstrate that STD-LLM exhibits strong generalization capabilities while maintaining competitive performance. Our work suggests that constructing a unified pre-trained large spatial-temporal model based on LLMs is a promising direction.
@@ Appendix A Limitations
Due to computational limitations, this paper did not further test whether LLM could yield better results. In theory, the recently emerged LLMs with 7 billion, 30 billion, or even more parameters should possess stronger capabilities for processing spatial-temporal data. Additionally, although this paper proposed a unified pre-training framework, we lack sufficient data to test whether the model, after being trained on a large-scale spatial-temporal dataset, would exhibit emergent phenomena similar to those of LLMs, leading to a substantial improvement in performance. These are some of the limitations of our research presented in this paper, and we hope that future work will be supported by adequate computational power and data to further explore these possibilities.
@@ C.1 Baselines
  • ASTGNN: An autoregressive forecasting model based on spatial-temporal attention and dynamic graph convolution. It improves the traditional multi-head attention module by introducing 1D convolution to process hidden local trend variations in the data. It also optimizes the conventional GCN by modeling spatial heterogeneity through additional embedding vectors.
@@
  • OFA: A time series model based on LLM (Large Language Model). It processes time series data by segmenting it into patches and converting it into tokens for LLM processing. It fine-tunes only the layer normalization and position encoding of LLM to retain its general capabilities.
  • STGLLM: A spatial-temporal forecasting model based on LLM. It enhances LLM’s understanding of spatial-temporal data using time embedding and other external information, as well as prompts. It also fine-tunes only the layer normalization and position encoding of LLM.
@@ C.4 Model Configurations
Spatial-Temporal embedding $d_{t}$ the dimension of time embedding 16
$d_{n}$ the dimension of topology-aware node embedding 16
$K$ the number of truncated eigenvectors 16
Spatial-Temporal Tokenizer $d_{T}$ the dimension of Temporal Tokenizer 64
Hypergraph Learning $M$ the number of region-level tokens 128
$d_{H}$ the dimension of hidden space in Hypergraph 128
LLM $d_{LLM}$ the dimension of the word embeddings in LLM 768
Layers the number of LLM layers utilized 3
@@ C.5 Parameter Efficacy
An important question to consider is whether the LLM significantly increases the computation cost. We have summarized the computation cost of main baselines and STD-LLM in Table 8. It can be observed from the table that, although our model has a large number of parameters, only 2.69% of them require training, and our model also has the fastest inference speed. The incorporation of LLM does indeed substantially increase the number of parameters, which leads to higher GPU memory usage. However, thanks to the relatively simple structure of LLM, all operations can be highly parallelized, thus not significantly impacting the training and inference time. Additionally, the regional-level tokens constructed by our hypergraph module also reduce the number of tokens, thereby enhancing computation speed.
@@
STD-LLM (ours) 64754907 1746520 2.69 5.53
Additionally, we have further tested the impact of the hypergraph module on the model’s inference efficiency. The results of the experiments are summarized in Table 9, we record the time cost to infer the entire test set (in seconds).
Table 9: Comparison of inference times with the batchsize=64.
Method PEMS03 PEMS04 PEMS07 PEMS08
w hypergraph 9.27 6.23 12.33 5.53
w/o hypergraph 19.973 11.50 55.32 7.31
From the results, we can observe that the hypergraph module significantly enhances the efficiency of the model, particularly on large graph structures. On the PEMS07 dataset with 883 nodes, the hypergraph module can improve inference efficiency by nearly five times.
@@ Appendix D Error Bars
All experiments were conducted three times, and the final metrics were taken as the average of the three trials. We present the standard deviations for STD-LLM in the forecasting and imputation experiments in Table 10 and Table 11.
Table 10: Standard deviations of STD-LLM in the forecasting experiment.
@@
MAE $14.89\pm 0.0636$ $18.31\pm 0.0321$ $19.71\pm 0.1153$ $13.85\pm 0.0011$
RMSE $25.38\pm 0.3181$ $30.32\pm 0.0458$ $32.99\pm 0.1374$ $23.43\pm 0.1123$
MAPE(%) $15.63\pm 0.2616$ $11.96\pm 0.0251$ $8.32\pm 0.0611$ $9.01\pm 0.0230$
Table 11: Standard deviations of STD-LLM in the imputation experiment on PEMS08.
Metric RM 30% RM 70% CM 30% CM 70%
MAE $12.14\pm 0.0321$ $14.38\pm 0.0871$ $19.36\pm 0.1352$ $22.18\pm 0.0793$
RMSE $19.44\pm 0.0400$ $22.92\pm 0.0700$ $34.53\pm 0.2150$ $39.26\pm 0.4179$
MAPE(%) $8.06\pm 0.0360$ $9.46\pm 0.1014$ $12.30\pm 0.0642$ $13.56\pm 0.1212$
Appendix E Detailed Definition for Constraint Loss
We decompose the constrain loss $\mathcal{L}_{C}$ in Equation 9 into $\mathcal{L}_{G}$ and $\mathcal{L}_{R}$:
$\mathcal{L}_{G}=-\sum_{m=1}^{M}\sum_{i,j(i\neq j)}{S}_{m,i}{S}_{m,j}A_{i,j},$ (11)
$\mathcal{L}_{R}=\log\pi(\text{Softmax}(\sum_{m=1}^{M}S_{m,:})|\bm{\alpha}),$
Firstly, as the attention score, $S_{m,:}$ satisfies the properties: $\sum_{i=1}^{N}S_{m,i}=1$.It is clear that the more edges there are, the smaller the minimum value of $L_{G}$ can be. Consider a complete graph of size P, where:
$\mathcal{L}_{G}$ $=-\sum_{m=1}^{M}\sum_{i,j(i\neq j)}{S}_{m,i}{S}_{m,j},$ (12)
$=-\sum_{m=1}^{M}\sum_{i}{S}_{m,i}(1-{S}_{m,i})=-M+\|{S}\odot{S}\|_{1}.$
$\odot$ means Hadamard Product. It is clear that the way to minimize $\mathcal{L}_{G}$ is to distribute the $S$ evenly, that is, $S_{m,i}=\frac{1}{P}$, at which point $\mathcal{L}_{G}=-\frac{M(P-1)}{P}$. When we shift our focus to a regular graph of size $N$, the smaller the $\mathcal{L}_{G}$, the more readily the hypergraph module can focus attention on larger complete subgraphs or subgraphs with stronger connectivity, thereby integrating the graph structure into the model. Of course, the premise of doing this is that the graph we are dealing with is a relatively sparse graph, containing many small complete subgraphs or subgraphs with strong connectivity. Otherwise, the hypergraph module might focus attention only on a few large complete subgraphs. If the graph we are facing is composed of just a few large complete subgraphs, then it is possible to achieve better results using prior knowledge.
Regarding $\mathcal{L}_{G}$, we intend for it to serve a guiding role during the training process and do not wish for the model to excessively reduce $\mathcal{L}_{G}$. Doing so could very likely lead to overfitting, which might result in extremely low attention being allocated to some nodes. Therefore, we introduce a regularization term $\mathcal{L}_{G}$. $\mathcal{L}_{R}$ is based on the Dirichlet distribution $\pi$, which is defined as follows:
$\Theta=\{\theta_{1},\theta_{2},\cdot,\theta_{N}\},\sum_{i}^{N}\theta_{i}=1,$ (13)
$\bm{\alpha}=\{a_{1},a_{2},\cdot,a_{N}\},$
$\pi(\Theta|\bm{\alpha})=\frac{\Gamma(\sum_{i}a_{i})}{\prod_{i}^{N}{\theta_{i}}^{a_{i}-1}}.$
The expected value of the Dirichlet distribution $E(\theta_{i})=\frac{a_{i}}{\sum_{i}a_{i}}$. Thus, we can set $\bm{\alpha}=\{1.05\}^{N}$(a small value greater than 1), which constrains the hypergraph module to aggregate each node equally. In practice, we believe that nodes of greater significance should receive more attention. Therefore, we incorporate the degree of the nodes: $\bm{\alpha}=\{1.05\}^{N}+c\times\text{Softmax}(\text{diag}(D))$. Here, $diag$ refers to the operation of extracting the diagonal elements of a matrix, and $c$ is a coefficient.
Appendix F Few-shot Performance on Pre-trained STD-LLM
In Section 4.4, we verified that the STD-LLM, which had not been pre-trained on spatial-temporal data, also possesses strong few-shot performance. So, would an STD-LLM that has been pre-trained on a specific spatial-temporal dataset exhibit even stronger few-shot performance? Can a pre-trained STD-LLM play a positive role in the training on other spatial-temporal datasets? To this end, we conduct some experiments to verify. We tested the few-shot performance and full dataset training performance of the STD-LLM pre-trained on PEMS07 for the PEMS08 and PEMS04 datasets, with the results summarized in Table 12.
Table 12: Few-shot on pre-trained STD-LLM.
Ratio PEMS04 PEMS08
MAE RMSE MAPE(%) MAE RMSE MAPE(%)
5% 22.11 35.24 14.76 17.05 27.16 10.92
10% 21.03 33.71 13.75 17.10 27.41 10.83
20% 20.21 32.99 13.18 16.01 26.12 10.56
100% 18.20 30.13 11.94 13.77 23.52 8.98
The results indicate that the few-shot performance of the pre-trained STD-LLM has seen a significant enhancement, achieving comparable performance to that of the non-pre-trained model using 20% of the data with just 5% of the data. Additionally, the performance of the model trained on the full dataset has also improved, demonstrating that STD-LLM can correctly transfer the knowledge acquired from pre-training to new tasks.
Appendix G Visualization
We have plotted some results for forecasting and imputation. It can be observed that STD-LLM can accurately follow the trend of data changes and also make corresponding adjustments to sudden changes in the data.
Figure 4: Traffic flow forecasting.
Figure 5: Traffic flow imputation. The green areas indicate the locations where data is missing and requires imputation.

### v2

@@ 4.1 Datasets
The experiments conducted on four traffic datasets (PEMS03, PEMS04, PEMS07, PEMS08) [44].
To assess the model’s imputation capability, we generated two types of missing data patterns, RM (random missing) and CM (spatial-temporal continuity missing), on the PEMS08 dataset, each with a 70% missing rate. Additionally, following the self-supervised training approach of CSDI  [35], we generated condition missing for training. Condition missing refers to the artificially created missing data, used to generate training samples.
The details of the four datasets and missing pattern are provided in supplementary material.
Table 1: Forecasting performance. There are two kinds of baselines, including six conventional deep learning models and three PLM-based models.
Model PEMS03 PEMS04 PEMS07 PEMS08
MAE RMSE MAPE MAE RMSE MAPE MAE RMSE MAPE MAE RMSE MAPE
LSTM 20.62 33.54 28.94% 26.81 40.74 22.33% 29.71 45.32 14.14% 22.19 33.59 18.74%
ASTGCN 17.85 29.88 17.65% 22.42 34.75 15.87% 25.98 39.65 11.84% 18.86 28.55 12.50%
AGCRN 15.98 28.25 15.23% 19.83 32.26 12.97% 22.37 36.55 9.12% 15.95 25.22 10.09%
SSTBAN 15.90 26.11 17.27% 18.88 31.10 12.56% 20.17 33.45 8.93% 14.39 24.19 10.06%
PDFormer 14.74 25.59 15.35% 18.31 29.97 12.10% 19.83 32.87 8.53% 13.58 23.51 9.05%
iTransformer 19.48 31.20 17.84% 22.56 35.21 16.29% 24.70 37.91 11.40% 20.05 31.90 11.99%
OFA 20.96 33.43 19.11% 27.37 42.99 17.97% 30.53 47.51 12.98% 21.89 34.63 13.30%
STGLLM 15.26 24.11 15.73% 20.00 32.11 13.69% 21.98 35.02 9.72% 15.53 24.74 10.15%
STLLM 17.25 27.25 22.96% 19.00 30.35 13.55% 21.48 34.07 10.20% 14.67 23.50 10.63%
STD-PLM 14.59 25.36 14.92% 18.16 30.21 11.89% 19.25 32.84 8.06% 13.31 23.19 8.84%
@@ 4.2 Experimental Settings
To comprehensively compare the performance of models, we selected nine representative baselines for the forecasting task, including Long Short-Term Memory network(LSTM) [45], ASTGCN [22], AGCRN [20], SSTBAN [40], PDFormer [24], iTransformer [46], OFA [37], STLLM [7] and STGLLM [6]. For the imputation task, we also chose four well-known baselines, which are Brits [28], E2GAN [29], mTAN [33], and PriSTI [36].
The experiments were conducted on NVIDIA A40 GPUs with a PyTorch version of 2.1.2. Using an AdaW optimizer with a learning rate of $1\times 10^{-3}$. The training epoch is 500 with a 50 epochs early stopping mechanism. We use GPT-2 [47] as the PLM and utilized only its initial three layers. STD-PLM employs the hyperparameters identified through tuning on the PEMS08 validation set for all datasets. The detail of experiments settings are depicted in supplementary material.
Table 2: Imputation performance on PEMS08.
Model RM 70% CM 70%
MAE RMSE MAPE MAE RMSE MAPE
BRITS 29.56 41.52 10.79% 38.33 53.60 13.90%
E2GAN 27.55 41.99 17.52% 31.58 50.24 19.17%
mTAN 21.23 33.98 12.89% 31.42 49.10 18.96%
PriSTI 25.54 36.00 16.73% 22.93 40.90 14.43%
STD-PLM 14.36 23.20 9.58% 22.69 39.66 13.82%
@@ 4.3 Overall Performance
We measure the model’s performance using three widely used metrics for regression tasks: Mean Absolute Error (MAE), Root Mean Square Error (RMSE), and Mean Absolute Percentage Error (MAPE). Table 1 shows the forecasting performances comparison between our proposed STD-PLM and baselines. Bold indicates the best results, while underline denotes the second-best results. From Table 1, we can observe that: 1) The accuracy of the time series models such as LSTM, iTransformer and OFA is not good enough compared to methods that can handle spatial-temporal correlations in data. This indicates that relying solely on temporal correlation is insufficient to achieve high-precision forecasting; 2) The results of PLM-based models OFA, STGLLM, and STLLM demonstrate the necessity of our design, as simply using the PLM to process spatial-temporal data does not lead to optimal performance. 3) STD-PLM achieve the best or second-best results on the all datasets.
Table 2 presents the imputation performances comparison between our proposed STD-PLM and baselines. Based on the comparison, we can observe that: 1)Our method achieves state-of-the-art performance. 2) STD-PLM exhibits a substantial improvement over other baseline models in the RM 70% case. This implies that STD-PLM can fully harness the imputation capabilities of the PLM for spatial-temporal data.
In summary, the experimental results demonstrate that STD-PLM has strong performance and can achieve high accuracy in both forecasting and imputation tasks.
@@ 4.4 Few-shot and Zero-shot Performance
To extensively evaluate the performance of PLM on spatial-temporal tasks, we conduct few-shot and zero-shot experiments. The few-shot experiments evaluate the forecasting performance of STD-PLM on the PEMS04 and PEMS08 datasets using only the first 5%, 10%, and 20% of the training samples, with results summarized in Table 4. The zero-shot experiments evaluate the forecasting performance of the trained model when applied to other datasets, with the results compiled in Table 4.
From Table 4, it can be observed that we only require 5% of the training samples to achieve performance comparable to that of LSTM trained on the full dataset. When the training samples are increased to 20%, the performance surpasses that of ASTGCN. This indicates that STD-PLM has excellent few-shot learning capabilities, which supports its application in scenarios where data is scarce. The results in Table 4 demonstrate that STD-PLM can exhibit acceptable performance when directly transferred to datasets with different temporal scopes and graph structures from the training set, without undergoing any training.
@@
MAE RMSE MAPE MAE RMSE MAPE
5% 27.73 42.12 20.62% 22.56 34.35 16.91%
10% 25.11 38.68 17.65% 19.72 31.35 13.27%
20% 21.16 34.03 13.92% 16.58 27.03 10.78%
@@
MAE RMSE MAPE
$\xrightarrow{}$08 29.52 45.63 22.92%
$\xrightarrow{}$04 25.32 37.80 25.32%
$\xrightarrow{}$03 23.72 36.94 41.59%
$\xrightarrow{}$07 34.72 52.66 20.31%
@@
{floatrow}
[…]
Method PEMS08 PEMS08 CM 70%
MAE RMSE MAPE MAE RMSE MAPE
w/o TT 13.81 23.52 9.03% 22.82 39.50 14.35%
w/o CLoss 13.53 23.69 8.97% 23.62 40.05 14.65%
w/o PLM 13.63 23.39 9.15% 24.05 41.15 15.34%
STD-PLM 13.31 23.19 8.84% 22.69 39.66 13.82%
[…]
Method PEMS03 PEMS07
Time Memory Time Memory
w SGA 7.40 8554 9.15 15020
w/o SGA 17.96 15366 52.82 29718
Table 5: Ablation study.
Table 6: inference consumption.
@@ 4.5 Ablation Study
We implement ablation study to evaluate the effectiveness of each component. The ablation experiments compared our proposed model with the following variants: 1) w/o TT, removing the temporal tokens; 2) w/o CLoss, removing the $\mathcal{L}_{C}$ in Eq.9; 3) w/o PLM, replacing the PLM with a transformer encoder of the same layer and hidden dimension.
We summerize the results of the ablation study in Table 6. Based on the results, we can identify the following: 1) Temporal Token and $\mathcal{L}_{C}$ can enhance the model’s forecasting and imputation performance with a relatively small computational cost. 2) Fine-tuning the PLM can achieve better results than a transformer encoder with the same parameter scale that updates all parameters during training. That means that the knowledge embedded within the PLM is useful for spatial-temporal tasks.
Additionally, we test the impact of the SGA on efficiency. The results of the experiments are summarized in Table 6, we record the time (s) and gpu memory (MiB) usage to infer the entire test set. From this, we can observe that the efficiency improvement brought by the SGA is substantial.
@@ 5 Conclusion
In this paper, we propose STD-PLM, a unified framework for spatial-temporal forecasting and imputation based on PLM. Through our explicitly designed spatial-temporal tokenizers and spatial-temporal embeddings, STD-PLM can effectively understand both spatial and temporal properties of spatial-temporal data. Additionally, we introduce a SGA module to significantly reduces computational costs by constructing region-level spatial tokens. Extensive experiments demonstrate that STD-PLM exhibits competitive performance. Our work suggests that constructing a unified pre-trained spatial-temporal model based on PLM is promising.
@@ Appendix A Limitations
Due to computational limitations, this paper did not further test whether PLM could yield better results. In theory, the recently emerged PLM with 7 billion, 30 billion, or even more parameters should possess stronger capabilities for processing spatial-temporal data. Additionally, although this paper proposed a unified pre-training framework, we lack sufficient data to test whether the model, after being trained on a large-scale spatial-temporal dataset, would exhibit emergent phenomena similar to those of PLM, leading to a substantial improvement in performance. These are some of the limitations of our research presented in this paper, and we hope that future work will be supported by adequate computational power and data to further explore these possibilities.
@@ C.1 Baselines
  • SSTBAN: An attention based long-term spatial-temporal prediction model using an autoencoder architecture. It designs a bottleneck attention that can effectively reduce the time complexity of the model’s attention process through some reference points. At the same time, the model also combines a mask autoencoder to effectively enhance the encoding ability of the model for spatial-temporal data through completion tasks.
@@
  • iTransformer: A transformative approach to time series forecasting that inverts the traditional Transformer model. By focusing on variate-centric embeddings and self-attention mechanisms, it adeptly captures multivariate correlations and learns robust nonlinear representations, achieving state-of-the-art results with enhanced efficiency and generalization.
  • OFA: A time series model based on PLM. It processes time series data by segmenting it into patches and converting it into tokens for PLM processing. It fine-tunes only the layer normalization and position encoding of PLM to retain its general capabilities.
  • STGLLM: A spatial-temporal forecasting model based on PLM. It enhances PLM’s understanding of spatial-temporal data using time embedding and other external information, as well as prompts. It also fine-tunes only the layer normalization and position encoding of PLM.
  • STLLM: A spatial-temporal prediction model based on PLM employs a special partial fine-tuning technique for the PLM. It fine-tunes only the Layer Normalization for the initial layers of the PLM, while for the subsequent layers, both the attention layers and Layer Normalization are fine-tuned simultaneously.
@@ C.4 Model Configurations
Spatial-Temporal embedding $d_{t}$ the dimension of time embedding 64
$d_{n}$ the dimension of topology-aware node embedding 64
$K$ the number of truncated eigenvectors 64
SGA $M$ the number of region-level tokens 128
$d_{H}$ the dimension of hidden space in SGA 128
PLM $d_{PLM}$ the dimension of the word embeddings in PLM 768
Layers the number of PLM layers utilized 3
@@ C.5 Parameter Efficacy
An important question to consider is whether the PLM significantly increases the computation cost. We have summarized the computation cost of main baselines and STD-PLM in Table 9. It can be observed from the table that, although our model has a large number of parameters, only 2.69% of them require training, and our model also has the fastest inference speed. The incorporation of PLM does indeed substantially increase the number of parameters, which leads to higher GPU memory usage. However, thanks to the relatively simple structure of PLM, all operations can be highly parallelized, thus not significantly impacting the training and inference time. Additionally, the regional-level tokens constructed by our SGA module also reduce the number of tokens, thereby enhancing computation speed.
@@
STLLM 82640908 42494476 51.42 24.42
STD-PLM (ours) 66136848 3128461 4.73 4.38
@@ Appendix D Error Bars
All experiments were conducted three times, and the final metrics were taken as the average of the three trials. We present the standard deviations for STD-PLM in the forecasting and imputation experiments in Table 11 and Table 11.
{floatrow}
[…]
@@
MAE $14.59\pm 0.0163$ $18.16\pm 0.0307$ $19.25\pm 0.1018$ $13.31\pm 0.0249$
RMSE $25.36\pm 0.1582$ $30.21\pm 0.0193$ $32.84\pm 0.0815$ $23.19\pm 0.0857$
MAPE(%) $14.92\pm 0.0727$ $11.89\pm 0.0429$ $8.06\pm 0.0383$ $8.84\pm 0.0214$
[…]
Metric RM 70% CM 70%
MAE $14.36\pm 0.0661$ $22.69\pm 0.0768$
RMSE $23.20\pm 0.2201$ $39.66\pm 0.5531$
MAPE(%) $9.58\pm 0.0846$ $13.82\pm 0.1213$
Table 10: Standard deviations of STD-PLM in the forecasting experiment.
Table 11: Standard deviations of STD-PLM in the imputation experiment on PEMS08.